Find the surface area of each pyramid to the nearest whole number?
11in. = slant height
12in. = perimeter

Answer by checkley71(8403)

what you have is 4 triangles with:
base=12/4=3in. & a height of 11in.
thus the area equation of the tringle is:
1/2Bh 1/2*3*11=33/2 sqin.
seeing as there are 4 sides then the surface area is:
4*1/2*3*11=66in.^3.